A moped license plate has 2 letters and then 4 numbers in it . How many plates can be made without duplicating, there are no plates with number 0 ?

Grade:11

2 Answers

Saurabh Koranglekar
5 years ago
26C4*9C4* 6!

simplify to get answer
Avik Denra
2 years ago
Total alphabets = 26
Total digits = 10
Total no. of plates = 26×26×10×10×10×10
                                = 6760000
Total plates with number zero = 26×26×1×1×1×1
                                                     = 676
Required answer = 6760000 - 676
Solve this for answer